# LADDER: Self-Improving LLMs Through Recursive Problem Decomposition

### v2

@@ 4.1 Llama 3B Experiments
Claude-3.5GPT-4oBase ModelLADDERo1-mini3030303040404040505050506060606070707070808080803232323242424242505050507373737380808080Score (%)
@@ 4.2 MIT integration Bee (LADDER)
For comparison, most human participants score between 3-6 points out of 20 (15-30%), with scores above 7-12 (35-73%) considered strong performances. With a 73% accuracy, our model meets the qualification threshold of 14/20. However, while LADDER achieved substantial gains, it still did not reach the performance of o1-mini (80%).
@@ 4.3 MIT Integration Bee (LADDER + TTRL)
After applying LADDER to the 7B base model, we further employed Test-Time Reinforcement Learning (TTRL) for 100 steps on questions that LADDER failed to answer correctly. Of the remaining incorrect responses, TTRL successfully solved 3 additional problems, increasing performance from 73% to 90% (Figure 6). This improvement places our model’s performance well above the typical qualification threshold of 73% for the MIT Integration Bee, setting a new state-of-the-art for mid-sized LLMs on this benchmark.
Deepseek R1LADDER pass@100o1LADDER + TTRL6060606070707070808080809090909010010010010078787878808080808080808090909090Score (%)
Figure 6: TTRL score improvement on 2025 MIT Integration Bee. Average across 3 runs for o1 and Deepseek R1.
The number of RL steps required to solve these additional questions varied, ranging from as few as 3 steps to as many as 30. This variation likely reflects the complexity of each integral, with more intricate problems requiring a greater number of solved variants to develop a useful learning trajectory. The questions that remained unsolved even after LADDER + TTRL are particularly complex for undergraduate students and are included in the appendix.
To assess the necessity of LADDER in TTRL’s effectiveness, we also applied TTRL directly to the base model. After 100 steps, the base model remained unable to correctly answer any of the questions that LADDER + TTRL successfully solved. Performance gains from TTRL only occurred when applied to the LADDER-trained model, suggesting that the RL protocol in LADDER provides essential exposure to the distribution of integrals, enabling TTRL to later learn further. Future research could explore whether TTRL alone, with sufficiently high compute constraints, could achieve comparable results.
@@ 6 Conclusion
After applying LADDER and TTRL, there were five questions on the 2025 MIT Integration Bee qualifying exam which continued to be answered incorrectly. The questions incorrectly answered are among the most complex questions on the exam, and for an undergraduate student are very difficult to solve.

### v3

@@ 4.1 Llama 3B Experiments
Claude-3.5GPT-4oBase ModelLADDERo13030303040404040505050506060606070707070808080803232323242424242505050507373737380808080Score (%)
@@ 4.2 MIT integration Bee (LADDER)
For comparison, most human participants score between 3-6 points out of 20 (15-30%), with scores above 7-12 (35-73%) considered strong performances. With a 73% accuracy, our model meets the qualification threshold of 14/20. However, while LADDER achieved substantial gains, it still did not reach the performance of o1 (80%).
@@ 4.3 MIT Integration Bee (LADDER + TTRL)
After applying LADDER to the 7B base model, we further employed Test-Time Reinforcement Learning (TTRL) for 100 steps on questions that LADDER failed to answer correctly. Of the remaining incorrect responses, TTRL successfully solved 3-4 additional problems, increasing performance from 73% to 90% (Figure 6). This improvement places our model’s performance well above the typical qualification threshold of 73% for the MIT Integration Bee, setting a new state-of-the-art for mid-sized LLMs on this benchmark.
LADDER pass@1LADDER pass@100LADDER + TTRL60606060707070708080808090909090100100100100o1 pass@1737373738080808090909090Score (%)
Figure 6: TTRL score improvement on 2025 MIT Integration Bee.
In Figure 6, we compare LADDER pass@100 with TTRL, demonstrating that TTRL is a more effective test-time scaling method than naively attempting a test question many times. TTRL’s application of recursive problem decomposition on the test questions themselves allows the initial LADDER model to correctly solve some of the hardest questions on the MIT Integration Bee qualifying examination. TTRL enables our relatively small 7B LADDER model to outperform OpenAI’s o1 model at pass@1.
The number of RL steps required to solve each test question previously incorrectly answered varied, ranging from as few as 3 steps to as many as 30. This variation likely reflects the complexity of each integral, with more complex problems requiring a greater number of solved variants to develop a useful learning trajectory. The questions that remained unsolved even after LADDER + TTRL are particularly complex for undergraduate students and are included in the appendix.
To assess the necessity of LADDER in TTRL’s effectiveness, we also applied TTRL directly to the base model (pre-LADDER). After 100 steps, the base model remained unable to correctly answer any of the questions that LADDER + TTRL successfully solved. Performance gains from TTRL only occurred when applied to the LADDER-trained model, suggesting that the RL protocol in LADDER provides essential exposure to the distribution of integrals, enabling TTRL to later learn further. Future research could explore whether TTRL alone, with sufficiently high compute constraints, could achieve comparable results.
@@ 6 Conclusion
After applying LADDER and TTRL, there were two questions on the 2025 MIT Integration Bee qualifying exam which continued to be answered incorrectly. The questions incorrectly answered are among the most complex questions on the exam, and for an undergraduate student are very difficult to solve.